A body dropped from a very large height , experiences resistance to its motion due to air and has avarying acceleration which decreases to zero in time t. Assume the displacement in this time and thevelocity acquired at the end of this time are same as that due uniform acceleration of cg (c is a constantless than 1). The body then travels with a uniform speed acquired at the end of time t. Find thedisplacement of the body in time 2t.


plzzz  solvee i will surely approve....

Grade:Upto college level

1 Answers

Pramod J AskiitiansExpert-IIT-B
12 years ago

Dear student,

velocity at the end of time t = v = 0 + cgt = cgt

displacement for the next time t = cgt*t = cgt^2

displacement for the first time t = s = 0 + 0.5cgt^2 = cgt^2/2

total displacement for the time 2t = (3/2)cgt^2

i hope this explains